You are thinking of purchasing a house. The house costs $400,000. You have $57,000 in cash that you can use as a down payment on the​ house, but you need to borrow the rest of the purchase price. The bank is offering a 30​-year mortgage that requires annual payments and has an interest rate of 5% per year. What will be your annual payment if you sign this​ mortgage?
Answer #1

Total borrowings=(400,000-57,000)=$343,000

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

343,000=Annuity[1-(1.05)^-30]/0.05

343,000=Annuity*15.37245103

Annuity=343,000/15.37245103

=$22312.64(Approx).